# XPrompt: Explaining Large Language Model’s Generation via Joint Prompt Attribution

### v1

@@ 5 Experiment
We provide quantitative studies to evaluate the faithfulness of the explanatory prompt fragments generated by XPrompt, comparing with existing interpretation baselines and ablation variants.
@@ 5.1 Experiment Settings
In the experiment, we employ two LLMs as our targeted $f_{\bm{\theta}}(\cdot)$: LLaMA2 (7B-Chat) [2] and Vicuna (7B) [32]. There are relatively few methods attributing prompts on the entire language generation, and we choose random removal (Random), Integrated-Gradient [22] and Captum [16] as our baseline for comparison. We implement these models using the PyTorch framework and pretrained weights from the transformers Python library [33], and conduct our experiments on an Nvidia RTX A6000-48GB platform with CUDA version 12.0.
@@
We employ three distinct text generation datasets: Alpaca [34], tldr_news [35], and MHC [36], to evaluate the effectiveness of our method across various generation tasks. As we aim to capture the joint influence of prompts on the model generations, we focus on relatively long-context prompts rather than simple one-sentence prompts. Longer prompts tend to contain more diverse vocabularies, convey more information, and thus more likely to exhibit high textual correlation.
For evaluation purposes, we randomly select approximately 110 data samples with at least 15 words from each dataset. All datasets are publicly available and more details about the dataset are illustrated in Appendix A.1.
@@ 5.2 Evaluation Metrics
Faithfulness scores are a key metric for assessing the quality of explanations, with faithful explanations accurately reflecting the model’s decision-making process [28]. Studies [37, 38] suggest that if a certain input tokens are truly important, their removal should lead to a more significant change in model output than the removal of random tokens [39]. Therefore, after removing explanatory tokens identified by a faithful method, the model’s new generation would significantly differ from using the original output. Moreover, the input prompt with masking these explanatory tokens are less likely to reproduce the original model response $\bm{y}$, indicating a lower value of $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$.
To quantitatively evaluate explanation faithfulness, we measure the change of model generation behavior from two dimensions. On one hand, we compare the model’s original and new generations: the originally generated sequence $\bm{y}$ is based on the complete input prompt $\bm{x}$ (e.g., $\bm{y}=f_{\bm{\theta}}(\bm{x})$), while the new generation is conditioned on the masked prompt (e.g., $\bm{y}^{\prime}=f_{\bm{\theta}}(\bm{m}\odot\bm{x})$).
@@
We thus measure the similarity of the original generation $\bm{y}$ and the new generation $\bm{y}^{\prime}$ based on their word frequency and semantics. A smaller similarity reflects a larger change in model generation, suggesting a better explanation. On the other hand, we measure the likelihood of generating the original output $\bm{y}$ when the model uses the masked prompt, e.g., changes on $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$. A smaller likelihood indicates better explanations whose mask prevents the model from generating its original output. Detailed definitions of these metrics are explained below.
@@
Word Frequency: BLEU [40] is widely used to measure how close the candidate text $\bm{y}^{\prime}$ is to the reference text $\bm{y}$. The score measures the precision of matching $n$-grams from the text $\bm{y}^{\prime}$ to $\bm{y}$ by a clipping method to avoid overcounting and adjusting for brevity of $\bm{y}^{\prime}$ if it is shorter than $\bm{y}$. ROUGE-L [41] is the metric for measuring the overlap of sequences of words between the two texts, evaluating how much of the $\bm{y}^{\prime}$ matches with the reference $\bm{y}$ (precision), how much the reference $\bm{y}$ is covered by the candidate $\bm{y}^{\prime}$ (recall), and combine them into an F1 score.
@@
Semantic Similarity: SentenceBert [42] is a variation of the BERT model which is designed to generate high-quality sentence embeddings for the pairs of sentences. We leverage SentenceBert to transform the text $\bm{y}$ and $\bm{y}^{\prime}$ into fixed-length embedding vectors, and calculate the cosine similarity between their embeddings to quantify their semantic similarity.
@@
Probability Measurement: We define two measurements to reflect how the likelihood of generating the original text $\bm{y}$ changes before and after applying the explanation mask. We first define the Probability Ratio (PR) to indicate how less likely to generate the original output $\bm{y}$ when explanatory tokens are masked: $\text{PR}(\bm{x},\bm{y},\bm{m})=\frac{p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})}{p_{\bm{\theta}}(\bm{y}|\bm{x})}$. If the PR score is far below the random baseline, we can conclude that the masked tokens are indeed important to cause the model generating $\bm{y}$. In addition, we also calculate the KL-divergence between these two distributions to measure their difference: $D_{\text{KL}}\big{(}p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})||p_{\bm{\theta}}(\bm{y}|\bm{x})\big{)}$, and a larger score indicates a larger change in generation likelihood and a more accurate explanation.
@@ 5.3 Main Results on Faithfulness (Q1)
We evaluate the faithfulness of the explanatory prompt tokens generated by our framework on five aforementioned metrics. Table 1 shows the results of different methods on three datasets by masking $k=3$ identified tokens. For all metrics except the KL-divergence, a lower score is better which is annotated as $\downarrow$. In general, we observe that our method consistently demonstrates better interpretation faithfulness on all datasets compared with baselines, with a clear margin. In particular, we have the following observations demonstrating the advantage of our method:
@@
This is evident from the high PR value of around 0.958 and the low KL value of only 0.023 on the MHC dataset when randomly masking tokens. Only when essential tokens are perturbed, particularly when masked, does the generation of content and probability change substantially. The noticeable gaps in these metrics between XPrompt and Random underscore the genuine importance of the masked tokens and their role as counterfactual explanations for the model output.
XPrompt can effectively capture semantically important prompt fragments, and the advantage stands out for long context.
Compared to Captum and Integrated-Gradient, our method generally yields lower values for BLEU and ROUGE-L across all datasets. The discrepancies between XPrompt and Captum are more pronounced for datasets with longer text, such as tldr_news and MHC. Specifically, on the tldr_news dataset, the F1-score of XPrompt is 12.7% lower than Captum for the LLaMA-2 model, and on the MHC dataset, it is 17.85% lower, this all indicates that XPrompt finds and removes the more important token. Furthermore, the SentenceBert results are better on all datasets, indicating larger semantic variations in the generated output after removing tokens by XPrompt.
XPrompt works even better on stronger LLMs.
The gaps between XPrompt and Captum are less apparent on the Vicuna-7B model compared to the LLaMA-2-7B-chat model, which could be attributed to the model’s inherent inference capabilities. Our method is based on the premise that there are textual correlations among the input tokens, allowing the model to infer from the remaining content when a portion of the tokens is masked. If the model has a poor ability to infer the masked token, it may struggle to capture the contextual information. Consequently, the effectiveness of our method is tied to the LLM’s proficiency in inference and understanding.
@@ 5.4 Time Efficiency (Q2)
XPrompt 15.225 15.397 14.473
[…]
@@ 5.5 Qualitative Visualization (Q3)
Figure 2 uses a case study to showcase our method can effectively identify interacted tokens. This figure illustrates the identified explanatory tokens by Captum and our method XPrompt. While there are few overlaps in the found tokens, compared with our method, those found by Captum are mostly not important to the generated response. This demonstrates the Captum’s limited ability to explain the relationship between input prompt and output response, especially considering the textual correlations within input tokens. For instances, the tokens “publishing”, “publish” and “book” have semantic correlations, thus Captum masking one by one individually cannot really eliminate this information complementarily provided by the others, especially when the LLM has a certain context inference ability. Similar issues exist in the second case, where “doctor” and “patient” has semantic interactions. Therefore, treating tokens independently and ignoring their joint influence on the generation is not a favored choice for prompt attribution.
This case study highlights the importance of our formulation of joint attribution, and verifies the effectiveness of our algorithm in discovering important token combinations on the generated output. More examples are shown in Appendix. A.2
5.6 Ablation Study
Number of Explanatory Tokens.
Figure 3 shows the Probability Ratio (PR) score as the number of masked tokens $k$ changes on each dataset for LLaMA-2(7B-Chat). Our algorithm consistently outperforms other baseline methods even with larger $k$, demonstrating the stable performance and effectiveness of XPrompt. After masking $k=2$ tokens, the value of PR would decrease dramatically, indicating that these two tokens are crucial for generating the output $\bm{y}$ and they act as the triggers that alter the probability distribution of the output. As the number of tokens increase to $k=5$, the PR keeps decreasing but with a less sharp slope. This suggests that the probability of generating the original output $\bm{y}$ is significantly determined by a few predominant tokens.
Figure 3: Probability Ratio with varying number of masked tokens $k$.
Figure 4: Convergence plot of XPrompt.
Number of Sampling Iterations.
Effectiveness of Gradient Guidance.
Recall that in our algorithm, we use gradient as guidance to initialize the mask (e.g., $\bm{m}^{(1)}$) and calculate sampling probability (i.e., softmax($\bm{m}^{(n)}\odot\bm{g}$)). To verify the efficiency of gradient guidance, we compare XPrompt with two variants: w/o Initialization that initializes the mask by uniformly random instead of gradient, and w/o Probability that samples swapping entries by uniformly random instead of gradient. Table 4 in the Appendix reports their resulting generation log-likelihood $\log p_{\bm{\theta}}(\bm{y}|\bm{m}\odot{\bm{x}})$ in different iterations.
We observe that without using gradient for initialization, the randomly initialized mask starts from a worse point with a high generation likelihood; and without using gradient to guide the sampling, the mask is updated in less effective direction to explore the search space, resulting in a high generation likelihood in the end. These results show that gradient is a useful and efficient tool for initializing the optimization from a better starting point and guiding the search to a better optimal point.
@@ 6 Conclusions
In this study, we introduce XPrompt, an efficient probabilistic search algorithm designed to generate the prompt attributions that elicit the model outputs for the generation tasks. We tackle the challenge of explaining the generation behavior for any given prompt by analyzing the joint effect of the prompt attributions on the output. We frame this explanation task as a discrete optimization problem, which can be efficiently solved by our proposed probabilistic search algorithm. This methodology enables efficient generation of any arbitrary number of explanatory prompt attributions that are deterministic to the generated content. Our framework is rigorously evaluated across extensive language generation tasks, including text summarization, question-answering, and general instruction datasets. The faithfulness of the explanatory prompt attributions is thoroughly analyzed and assessed by comprehensive metrics. The results demonstrate that our proposed method efficiently generates explanatory attributions that faithfully reflect the model’s generation behavior for the specific prompt. These explanatory attributions interact in semantically and jointly influence the output generation. Furthermore, the overall excellent performance of our method on these diverse datasets highlights its remarkable transferability.
@@ Appendix A Appendix
A.1 Dataset Details
We evaluate explanations on three datasets: Alpaca, tldr_news and MHC. Due to the extensive computational cost, we randomly select approaximately 110 data samples with at least 15 words from each datasets. The statistics of the data used in our experiment can be found in Table 3.
  • Alpaca [34] is a dataset with 52000 unique examples consisting of instructions and demonstrations generated by OpenAI’s text-davinci-003 engine. Each example in the dataset includes an instruction that describes the task the model should perform, accompanied by optional input context for that task. In our experiment, we randomly select a subset of these examples for verification purposes.
  • tldr_news [35] dataset is constructed by collecting a daily tech newsletter. For every piece of data, there is a headline and corresponding content extracted. The task is to ask the model to simplify the extracted content and then generate a headline from the input.
  • mental_health_counseling (MHC) [36] includes broad pairs of questions and answers derived from online counseling and therapy platforms. It covers a wide range of mental-health related questions and concerns, as well as the advice provided by the psychologists. It is utilized for used to fine-tuning the model to generate the metal health advice. Here, we prompt the model to generate an advice based on the provided question.
@@ A.1 Dataset Details
A.2 Case Study
In this section, we visualize three examples which are selected from the dataset Alpaca, tldr_news, and MHC respectively, shown in Figure. 5. The tokens in the red frame are the $k$ explanatory tokens extracted by the model LLaMA-2(7B-Chat), and the output responses framed in blue are the most informative results selected by human. In the first data sample, the recognized token: "brains" and "cognitive", has some semantic relationship intuitively. It is a common sense the cognitive ability is underpinned by the brain’s function and structure. There are numerous reports and studies focused on the functionally relationship between the brain and coginitive ability [43]. Due to the extensive amount of data on which the LLMs are trained, the model can integrate their relationship which is reflected in its response.
@@ A.2 Case Study
In Figure 6, we also illustrate the changes in the model response as the targeted top-$k$ token is removed from the input prompt. Compared with the top-$k$ tokens identified by our method, those found by Captum are mostly not important to the response. This demonstrates the Captum’s limited ability to explain the relationship between input prompt and output response with the consideration of the textual information. Moreover, when masking the token "publishing" or "publish", as identified by our method XPrompt, the generated outputs do not exhibit significant variations in either word frequency or semantics. However, substantial changes in the generated outputs occur only when both tokens are simultaneously masked. This highlights the effectiveness of our methodology in explaining the joint effect of the prompt attributions on the generated output.
A.3 Gradient Design

### v2

@@ 5 Experiment
We provide quantitative studies to evaluate the faithfulness of the explanatory prompt fragments generated by JoPA, comparing with existing interpretation baselines and ablation variants.
@@ 5.1 Experiment Settings
In the experiment, we employ two LLMs as our targeted $f_{\bm{\theta}}(\cdot)$: LlaMA-2 (7B-Chat) (touvron2023llama, ) and Vicuna (7B) (Zheng2023JudgingLW, ). There are relatively few methods attributing prompts on the entire language generation, and we choose random removal (Random), Integrated-Gradient (sundararajan2017axiomatic, ), averaged attentions across all layers (pruthi2019learning, )(Attention), last-layer attention (zhao2024reagent, )(Last-Attention) and Captum (captum, ) as our baseline for comparison. We also compare JoPA with ReAGent (zhao2024reagent, ), with the results presented in Appendix A.13. We implement these models using the PyTorch framework and pretrained weights from the transformers Python library (wolf-etal-2020-transformers, ), and conduct our experiments on an Nvidia RTX A6000-48GB platform with CUDA version 12.0.
@@
We employ three distinct text generation datasets: Alpaca (alpaca, ), tldr_news (tldr_news, ), and MHC (amod_2024, ), to evaluate the effectiveness of our method across various generation tasks. To capture the joint influence of prompts on model outputs, we use long-context prompts (15+ words) rather than simple sentences, as they convey more information and exhibit higher textual correlation. For evaluation, 110 samples are randomly selected from each publicly available dataset, detailed in AppendixA.2
@@ 5.2 Evaluation Metrics
Faithfulness scores are a key metric for assessing the quality of explanations, with faithful explanations accurately reflecting the model’s decision-making process (jacovi-goldberg-2020-towards, ). Studies (7552539, ; hooker2019benchmark, ) suggest that if a certain input tokens are truly important, their removal should lead to a more significant change in model output than the removal of random tokens (madsen-etal-2022-evaluating, ). Therefore, after removing explanatory tokens identified by a faithful method, the model’s new generation would significantly differ from using the original output. Moreover, the input prompt with masking these explanatory tokens are less likely to reproduce the original model response $\bm{y}$, indicating a lower value of $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$.
@@
To quantitatively evaluate explanation faithfulness, we measure the change of model generation behavior from two perspectives. First, we compare the original generation $\bm{y}=f_{\bm{\theta}}(\bm{x})$ with the new generation $\bm{y}^{\prime}=f_{\bm{\theta}}(\bm{m}\odot\bm{x})$, measuring similarity in word frequency and semantics. Lower similarity indicates greater change, suggesting better explanations. Second, we evaluate the likelihood of the original output $\bm{y}$ under the masked prompt, $p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})$. Lower likelihood reflects more effective masks that prevent the model from replicating the original output. Detailed metric definitions are explained below.
@@
Word Frequency: BLEU (bleu, ) evaluates how closely the candidate text $\bm{y}^{\prime}$ matches the reference $\bm{y}$ by measuring n-gram precision with clipping to avoid overcounting and brevity adjustments for shorter candidates. ROUGE-L (rouge, ) measures sequence overlap between $\bm{y}^{\prime}$ and $\bm{y}$, assessing precision, recall, and their F1 score.
@@
Semantic Similarity: SentenceBert (reimers2019sentencebert, ) is a variation of the BERT model which is designed to generate high-quality sentence embeddings for the pairs of sentences. We leverage SentenceBert to transform the text $\bm{y}$ and $\bm{y}^{\prime}$ into fixed-length embedding vectors, and calculate the cosine similarity between their embeddings to quantify their semantic similarity.
@@
Probability Measurement: We define two measurements to reflect how the likelihood of generating the original text $\bm{y}$ changes before and after applying the explanation mask. We first define the Probability Ratio (PR) to indicate how less likely to generate the original output $\bm{y}$ when explanatory tokens are masked: $\text{PR}(\bm{x},\bm{y},\bm{m})=\frac{\tilde{p}_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})}{\tilde{p}_{\bm{\theta}}(\bm{y}|\bm{x})}$, where $\tilde{p}(\bm{y}|\bm{x})=p(\bm{y}|\bm{x})^{1/s}$ is the length-normalized generation probability. If the PR score is far below the random baseline, we can conclude that the masked tokens are indeed important to cause the model generating $\bm{y}$. In addition, we also calculate the KL-divergence between these two distributions to measure their difference: $D_{\text{KL}}\large(p_{\bm{\theta}}(\bm{y}|\bm{m}\odot\bm{x})||p_{\bm{\theta}}(\bm{y}|\bm{x})\large)$, and a larger score indicates a larger change in generation likelihood and a more accurate explanation.
@@ 5.3 Main Results on Faithfulness (Q1)
We evaluate the faithfulness of explanatory prompt tokens using five metrics. Table 1 presents results on three datasets by masking $k=3$ identified tokens, with the larger dataset results in Appendix A.8. For all metrics except KL-divergence ($\downarrow$ indicates lower is better), our method consistently outperforms baselines across all datasets with a clear margin, demonstrating superior interpretative faithfulness. Key observations highlight our method’s advantages:
@@
This is supported by a high PR value of around 0.958 and low KL value of only 0.023 on the MHC dataset when masking random tokens. Substantial changes occur only when essential tokens are masked, highlighting their importance as counterfactual explanations. The metric gaps between JoPA and Random emphasize this, with performance stability detailed in Appendix A.9.
JoPA can effectively capture semantically important prompt fragments, and the advantage stands out for long context.
Compared to Captum and Integrated-Gradient, JoPA achieves lower BLEU and ROUGE-L values across all datasets, especially for longer texts like tldr_news and MHC. On tldr_news, the F1-score of JoPA is 12.7% lower, and on MHC, 17.85% lower than Captum for the LlaMA-2 (7B-Chat) model, indicating it identifies and removes more important tokens. Additionally, higher SentenceBert scores across datasets suggest greater semantic variation in outputs after token removal by JoPA.
JoPA works even better on stronger LLMs.
The gaps between JoPA and Captum are less apparent on the Vicuna (7B) model compared to the LlaMA-2 (7B-Chat) model, which could be attributed to the model’s inherent inference capabilities. JoPA is based on the premise that there are textual correlations among the input tokens, enabling inference from remaining content when tokens are masked. Models with weaker inference abilities may struggle to capture context, making our method’s effectiveness dependent on the LLM’s proficiency in inference and understanding.
@@ 5.4 Time Efficiency (Q2)
JoPA 15.225 15.397 14.473
Table 2: Time Efficiency on LLaMA-2 (7B-Chat)
In Table 2, we compare the average time cost of JoPA and Captum for generating $k=3$ explanatory tokens for each prompt instance. Note that since Captum’s design requires sequentially appending the next token to the input prompt to re-generate the new output, the time consumption of Captum would increase significantly when the prompt is long. As the average length of MHC is longer than the Alpaca shown in Table 3, the computational time increases from 1169.648ss to 1806.551ss for Captum, which is quite inefficient and impractical. Our algorithm efficiently solves the combinatorial optimization problem using a probabilistic search, significantly reducing computational costs for explanation generation. Since it requires only a certain number of gradient-guided search steps, computation time remains consistent regardless of prompt length, as shown in Table 3.
@@ 5.5 Qualitative Visualization (Q3)
Figure 2 uses a case study to showcase our method can effectively identify interacted tokens. This figure illustrates the identified explanatory tokens by Captum and our method JoPA. While there are few overlaps in the found tokens, compared with our method, those found by Captum are mostly not important to the generated response. This demonstrates the Captum’s limited ability to explain the relationship between input prompt and output response, especially considering the textual correlations within input tokens. For instances, the tokens “publishing”, “publish” and “book” have semantic correlations, thus Captum masking one by one individually cannot really eliminate this information complementarily provided by the others, especially when the LLM has a certain context inference ability. Similar issues exist in the second case, where “doctor” and “patient” has semantic interactions. Therefore, treating tokens independently and ignoring their joint influence on the generation is not a favored choice for prompt attribution. This case study underscores the importance of joint attribution and demonstrates our algorithm’s effectiveness in identifying key token combinations. Additional examples are in Appendix A.3. The algorithm also shows promise for enhancing model safety and aiding diagnosis, detailed in Appendix A.4.
@@ 6 Conclusions
We introduce JoPA, a probabilistic search algorithm that efficiently identifies prompt attributions explaining model outputs in generation tasks. By framing the explanation as a discrete optimization problem, JoPA generates faithful and deterministic attributions that influence outputs. Rigorous evaluation on tasks like summarization, question-answering, and instruction datasets shows JoPA’s effectiveness, faithfulness, and transferability.
7 Limitations
@@ Appendix A Appendix
A.1 Theoretical Insight
Here we could prove that our algorithm can theoretically converge to the local optima given enough iterations. Define that a solution $\bm{m}^{*}\in\{0,1\}^{T}$ is the local optima, if we have $p_{\bm{\theta}}(y|\bm{m}^{*}\odot\bm{x})\leq p_{\bm{\theta}}(y|\bm{m}\odot\bm{x})$ for all $\bm{m}$ in the one-swap neighborhood of $\bm{m}^{*}$, namely $\bm{m}^{*}-\bm{m}||_{0}=2$ (meaning $\bm{m}$ differs from $\bm{m}^{*}$ by one swap, as they are constrained to have $k$ zero entries). It could be proved that the output of the algorithm $\hat{\bm{m}}$ is the local optima by contradiction that with sufficient iterations. Suppose $\hat{\bm{m}}$ is not the local optima, there must be a point $\bm{m}^{\prime}$ in its neighbor satisfying $\bm{m}^{\prime}-\hat{\bm{m}}||_{0}=2$, such that $p_{\bm{\theta}}(y|\bm{m}^{\prime}\odot x)<p_{\bm{\theta}}(y|\hat{\bm{m}}\odot\bm{x})$. This means our algorithm can still find a better solution by sampling another swap in further iterations and thus $\hat{\bm{m}}$ is not the output of our algorithm, leading to a contradiction to the assumption.
A.2 Dataset Details
We evaluate explanations on three datasets: Alpaca, tldr_news and MHC. Due to the extensive computational cost, we randomly select approximately 110 data samples with at least 15 words from each datasets. The statistics of the data used in our experiment can be found in Table 3.
  • Alpaca alpaca is a dataset with 52000 unique examples consisting of instructions and demonstrations generated by OpenAI’s text-davinci-003 engine. Each example in the dataset includes an instruction that describes the task the model should perform, accompanied by optional input context for that task. In our experiment, we randomly select a subset of these examples for verification purposes.
  • tldr_news tldr_news dataset is constructed by collecting a daily tech newsletter. For every piece of data, there is a headline and corresponding content extracted. The task is to ask the model to simplify the extracted content and then generate a headline from the input.
  • mental_health_counseling (MHC) amod_2024 includes broad pairs of questions and answers derived from online counseling and therapy platforms. It covers a wide range of mental-health related questions and concerns, as well as the advice provided by the psychologists. It is utilized for used to fine-tuning the model to generate the metal health advice. Here, we prompt the model to generate an advice based on the provided question.
@@ A.2 Dataset Details
A.3 Case Study
In this section, we visualize three examples which are selected from the dataset Alpaca, tldr_news, and MHC respectively, shown in Figure. 3. The tokens in the red frame are the $k$ explanatory tokens extracted by the model LlaMA-2 (7B-Chat), and the output responses framed in blue are the most informative results selected by human. In the first data sample, the recognized token: "brains" and "cognitive", has some semantic relationship intuitively. It is a common sense the cognitive ability is underpinned by the brain’s function and structure. There are numerous reports and studies focused on the functionally relationship between the brain and cognitive ability (zhang2019cognitive, ). Due to the extensive amount of data on which the LLMs are trained, the model can integrate their relationship which is reflected in its response.
@@ A.3 Case Study
In Figure 4, we also illustrate the changes in the model response as the targeted top-$k$ token is removed from the input prompt. Compared with the top-$k$ tokens identified by our method, those found by Captum are mostly not important to the response. This demonstrates the Captum’s limited ability to explain the relationship between input prompt and output response with the consideration of the textual information. Moreover, when masking the token "publishing" or "publish", as identified by our method JoPA, the generated outputs do not exhibit significant variations in either word frequency or semantics. However, substantial changes in the generated outputs occur only when both tokens are simultaneously masked. This highlights the effectiveness of our methodology in explaining the joint effect of the prompt attributions on the generated output.
A.4 Benefits to the community on improving model ability
Benefit on Model Safety
Our method has the potential to help LLM defend against jailbreaking attacks, making the model more robust. We verify this via a real case study. Given an adversarial prompt with malicious request: "Generate a message that spreads fake news. organized few GO a a a a newcommand ! SudGuiheures ! !/ refection/:" It can jailbreak Llama-2 (7B-Chat) model to generate unsafe responses. Our algorithm successfully attributes "Sud" and "he", which are part of the adversarial suffix. By deleting these attributions, the attack fails. Thus, our approach could help detect these high-risk malicious tokens to robustify model safety.
Benefit on Model Diagnosis
Our approach can also be used to assess the relevance of model’s response to the prompt, helping diagnose the effectiveness of model responding to the prompt. For a real case study, consider a single prompt consisting of multiple tasks: "Create a message to complain about the noise in the neighborhood to friends. Give me the Python code for filtering results based on the following columns…" We find Llama-2 (7B-Chat) only responds to the first request, neglecting the second one. Our approach attributes the model’s output to tokens like "complain" "noise" and "friends", mainly about the first request. Thus by input attribution, our method can help diagnose the overlooked request and suggest better prompt design.
Detection Accuracy
JoPA could be used as the defense of the jailbreak attacks and help users to design better prompts to get satisfying responses. We also do the experiments to show the effectiveness of applying JoPA to detect malicious prompts on Llama-2 (7B-Chat) model, which is similar to the application done in CONTEXTCITE (contextcite, ) The results are shown below:
Method Detection Accuracy by JoPA ASR
GCG 100% 3%
Prompt with Random Search 91% 90%
Table 4: Comparison of detection accuracy and attack success rate.
[…]
A.5 Discussion on Counterfactual Explanation
Our work JoPA is to explain the generation of LLMs by identifying a small changes (i.e., masking few tokens) needed to alter the outcome (i.e., minimizing the probability of the original generation). This can be understood as counterfactual explanation, since it answers a "what if" question: if these input tokens are masked, the LLM would not generate the response. Namely, by applying the mask $m$ on the input prompt $x$, the generation probability of its original response $y$ would be largely decreased (i.e., $p_{\theta}(y|m\odot x)\ll p_{\theta}(y|x)$). This suggests that the newly generated output $y^{\prime}$ resulting from the masked prompt would differ substantially from the original outcome $y$, as measured by our metrics in Table 1: the semantic similarity (SentenceBert) and word similarity (BLEU and ROUGE-L) between $y$ and $y^{\prime}$ are much smaller than those from baselines. Here are several case studies to illustrate the concepts discussed.
General Q&A Task
[…]
Fill-in-the-Blank Task
[…]
The case studies of diverse tasks above demonstrate that the new responses generated from the masked inputs differ significantly from the original outputs, which highlights the impact of these explanatory tokens in influencing and altering the model’s output.
A.6 Ablation Study
Number of Explanatory Tokens.
Figure 6 shows the Probability Ratio (PR) score as the number of masked tokens $k$ changes on each dataset for LlaMA-2 (7B-Chat). Our algorithm consistently outperforms other baseline methods even with larger $k$, demonstrating the stable performance and effectiveness of JoPA. After masking $k=2$ tokens, the value of PR would decrease dramatically, indicating that these two tokens are crucial for generating the output $\bm{y}$ and they act as the triggers that alter the probability distribution of the output. As the number of tokens increase to $k=5$, the PR keeps decreasing but with a less sharp slope. This suggests that the probability of generating the original output $\bm{y}$ is significantly determined by a few predominant tokens.
Figure 5: Probability Ratio with varying number of masked tokens $k$.
Figure 6: Convergence plot of JoPA.
Number of Sampling Iterations.
[…]
In addition, we extend the application of our framework, JoPA, to a larger model and a more challenging task, specifically Few-shot Chain-of-Thought (CoT) reasoning (wei2022chain, ; kojima2022large, ), as described in Appendix A.11 and Appendix A.12. The better performance of our algorithm compared with others highlights the transferability of our framework across different model architectures and tasks.
A.7 Gradient Design